# Autoencoders for At-Source Data Reduction and Anomaly Detection in High Energy Particle Detectors

### v1

@@ 4 Anomaly Detection
A further capability of the on-detector encoder of a VAE is to detect anomalies at the edge using the track’s latent space encoding. The data-driven training of the VAE allows the model to learn the underlying data distribution and detect outlier events without the use of a signal model, enabling agnostic detection of unusual detector activity ranging from detector malfunctions [21, 22] to potential new physics. Furthermore, the ability to detect anomalies in real-time through a VAE-based readout scheme could allow the readout to dynamically decide track-by-track whether to compress the output or transmit it in full uncompressed precision for further study. Since outlier events are often the most challenging to accurately decompress, transmitting these events at full precision off-detector could help mitigate precision loss from this scheme.
To test the anomaly detection capability of the VAE described in Section 3, synthetic pixel events are constructed to represent various detector anomalies, starting from the original smart pixel tracks in the test set. Three different kinds of detector anomalies are considered: sensors with a single dead pixel, sensors with a single noisy pixel, or sensors with a dead row of pixels. Individual dead pixels are randomly chosen from the brighter half of pixels with nonzero charge. Dead pixels rows are randomly chosen from pixel rows with nonzero charge and their charge is set to a value of zero. The charge of a loud pixel in the loud pixel anomaly is 9000, determined by the 99th percentile of the max pixel value per track across the entire dataset. Figure 4 displays these anomalies as they appear in the pixel sensor, alongside a standard smart pixel track for reference. These synthetic anomalous tracks undergo the same input processing as the standard tracks in both the low- and high-precision scenarios.
@@
The encoder of the same VAE described in the previous section is evaluated over test sets of approximately 55,000 events for each of the three simulated anomalies, and the clipped KLD scores are calculated for each event for use as an anomaly classifier score. Figure 5 displays the anomaly score distributions of the standard track and anomaly test sets. Each of the anomalies is reconstructed with a higher mean anomaly score compared to the standard tracks, with a tail to higher values indicating a higher likelihood of incompatibility with the background. The VAE is least sensitive to dead pixel anomalies in the low-precision scenario. The anomaly detection capability was found to be anticorrelated with the reconstruction quality, indicating an area in need of further study and optimization for a realistic front-end scheme. Table 2 shows the achievable efficiency for identifying sensor anomalies for a fixed background track rejection of 90%.
@@
ML at any stage of collider DAQ systems must be implemented carefully to avoid alteration or corruption of the raw data. While VAE-based sensor data compression can mitigate high data rates during ordinary operation with minimal artifacts introduced in reconstruction, any indication of anomalous behavior in the sensors would ideally be preserved in original form. Thus, the compression and anomaly detection capabilities shown here could be used in tandem in a front-end readout design so that the events with the highest anomaly score are flagged and sent off the detector with full precision. With 1% of tracks with the highest anomaly score being transmitted uncompressed, this scheme could still achieve a data rate reduction of 16.4% of the original size while preserving the most unusual detector behaviors without compressed transmission for offline study.
@@ 5 Conclusions
A VAE capable of performing both data compression and anomaly detection in on-detector silicon pixel readout is presented. With an on-chip encoder and off-detector decompression, a VAE-based readout approach can reduce the off-detector data rate to 15.4% of its original size for 4-bit charge values, while providing a faithful reconstruction of the original pixel charge distribution. Further, information from the VAE latent space can enable anomaly detection with potential for real-time monitoring of pixel data quality. Both of these capabilities can be achieved with a latency less than 25 ns, better performance, and similar or fewer resources than a comparable on-chip track filtering approach.
Future work will seek to further test the fidelity of the decompressed data, specifically working to further reduce the reconstruction error and to study more complex reconstruction tasks such as feature regression or pattern recognition. The generality of this approach can also be leveraged to consider the integration of front-end VAEs into a variety of other detector technologies. A key step will be the assessment of potential hardware targets for these models and ensuring their viability for deployment in the extreme collider environment.

### v2

@@ 4 Anomaly Detection
While VAE-based data compression is a traditional application, this result extends to the novel capability of the encoder to detect anomalies at the edge using the track’s latent space encoding. The data-driven training of the VAE allows the model to learn the underlying data distribution and detect outlier events without the use of a signal model, enabling agnostic detection of unusual detector activity ranging from detector malfunctions [21] to potential new physics. Furthermore, the ability to detect anomalies in real-time through a VAE-based readout scheme could allow the readout to dynamically decide track-by-track whether to compress the output or transmit it in full uncompressed precision for further study. Since outlier events are often the most challenging for the VAE to accurately reconstruct, transmitting these events at full precision off-detector could help mitigate the loss of key information of anomalous pixel events.
To test the anomaly detection capability of the VAE described in Section 3, synthetic pixel events are constructed to represent various detector anomalies, starting from the original smart pixel tracks in the test set. Three different kinds of detector anomalies are considered: sensors with a single dead pixel, sensors with a single noisy pixel, or sensors with a dead row of pixels. These anomalies represent possible detector malfunctions that are important to identify and flag. Individual dead pixels are randomly chosen from the brighter half of pixels with nonzero charge. Dead pixels rows are randomly chosen from pixel rows with nonzero charge and their charge is set to a value of zero. The charge of a loud pixel in the loud pixel anomaly is 9000, determined by the 99th percentile of the max pixel value per track across the entire dataset. Figure 4 displays these anomalies as they appear in the pixel sensor, alongside a standard smart pixel track for reference. These synthetic anomalous tracks undergo the same input processing as the standard tracks in both the low- and high-precision scenarios.
@@
The encoder of the same VAE described in the previous section is evaluated over test sets of approximately 55,000 events for each of the three simulated anomalies, and the clipped KLD scores are calculated for each event for use as an anomaly classifier score. Figure 5 displays the anomaly score distributions of the standard track and anomaly test sets. Each of the anomalies is reconstructed with a higher mean anomaly score compared to the standard tracks, with a tail to higher values indicating a higher likelihood of incompatibility with the background. The anomaly detection capability was found to be anticorrelated with the reconstruction quality, indicating an area in need of further study and optimization for a realistic front-end scheme. Table 2 shows the achievable efficiency for identifying sensor anomalies for a fixed background track rejection of 90%.
@@
ML at any stage of collider DAQ systems must be implemented carefully to avoid alteration or corruption of the raw data. While VAE-based sensor data compression can mitigate high data rates during ordinary operation with minimal artifacts introduced in reconstruction, any indication of anomalous behavior in the sensors would ideally be preserved in original form. Thus, the compression and anomaly detection capabilities shown here could be used in tandem in a front-end readout design so that the events with the highest anomaly score are flagged and sent off the detector with full precision. With 1% of tracks with the highest anomaly score being transmitted uncompressed, this scheme could still achieve a data rate reduction of 16.4% of the original size while preserving the most unusual detector behaviors at full precision for offline study.
@@ 5 Conclusions
A VAE capable of performing both data compression and anomaly detection in on-detector silicon pixel readout is presented. With an on-chip encoder and off-detector decompression, a VAE-based readout approach can reduce the off-detector data rate to 15.4% of its original size for 4-bit charge values, while providing a faithful reconstruction of the original pixel charge distribution. Further, information from the VAE latent space can enable anomaly detection with potential for real-time monitoring of pixel data quality. Both of these capabilities can be achieved with a latency less than 25 ns, better performance, and similar or fewer resources than a comparable on-chip track filtering approach. This results opens the door to intelligent inference in front-end readout of modern collider experiments, with potential to dramatically improve the real-time filtering systems required to manage collider data rates, and via a generic architecture that can be applied to a wide variety of detector subsystems.
Future work will seek to further test the fidelity of the decompressed data, specifically working to further reduce the reconstruction error and to study more complex reconstruction tasks such as feature regression or pattern recognition. The resource constraints in a real detector deployment will be strongly dependent on the power requirements of this system, requiring careful evaluation of on-chip power needs and tailoring to the specific application. Additional comparison to traditional on-chip data compression approaches [22, 23, 24] will be essential motivation for the broad introduction of ML-based readout methodologies. The generality of this approach can also be leveraged to consider the integration of front-end VAEs into a variety of other detector technologies, such as for waveform analysis of calorimeter or drift chamber signals. A key step will be the assessment of potential hardware targets for these models and ensuring their viability for deployment in the extreme collider environment.